OpenStudy (anonymous):

what is x^2 - 18x + 82 factored?

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

i bet it is really \[x^2-18x+81\] right? not \(82\)

OpenStudy (anonymous):

the thing is i'm doing radicals so I had to move the one to the other side, so now i'm kind of stuck on what to do

OpenStudy (misty1212):

if it is what i wrote, then it is a perfect square, namely \[(x-9)^2\]

OpenStudy (misty1212):

oh dear, there is no 'other side" in your question, because you do not have an equal sign there

OpenStudy (anonymous):

Okay lets start over, maybe you could help me. Can i send you the question I am doing?

OpenStudy (misty1212):

is it \[x^2-18x+82=0\]?

OpenStudy (misty1212):

sure go ahead

OpenStudy (anonymous):

No lol here just take a look When solving a radical equation, John and Tim came to two different conclusions. John found a solution, while Tim's solution did not work in the equation. Create and justify two situations: one situation where John is correct and a separate situation where Tim is correct.

OpenStudy (misty1212):

hmm that is a different question entirely

OpenStudy (misty1212):

a radical equation could look like \[\sqrt{x-1}+3=5\]

OpenStudy (misty1212):

you would solve that by writing \[\sqrt{x-1}+3=5\\ \sqrt{x-1}=2\\ x-1=4\\ x=5\]

OpenStudy (anonymous):

Yes exactly, I already had one written, but then I got to x = x^2 - 18x +82 and got totally confused.

OpenStudy (anonymous):

you're suppose to square both sides?

OpenStudy (here_to_help15):

okay, sometimes when you solve an equation that has a radical sign in it, by squaring it, you end up introducing what is called an extraneous solution. An extraneous solution is not a mistake, but rather a solution that doesn't work in the original equation. There's a video here that might help you: https://www.khanacademy.org/math/algebra/exponent-equations/radical_equations/v/extraneous-solutions-to-radical-equations

OpenStudy (anonymous):

Okay, i see what you did

OpenStudy (misty1212):

yes squared both sides

OpenStudy (anonymous):

@Here_to_Help15 I know, I've seen that already, I'm just stuck on finding the right problem that will give me a extraneous and non-extraneous solution

OpenStudy (misty1212):

but if you had \[\sqrt{x-1}+5=3\] then \[\sqrt{x-1}=-2\] and if you square you get \[x-1=4\] so \[x=5\] but \(5\) is not really as solution, you can check it

OpenStudy (misty1212):

the first one i wrote has a real solution, the second one has an "extraneous" solution

OpenStudy (anonymous):

It has to have both, see this was my equation \[\sqrt{x-1} = x-9\]

OpenStudy (anonymous):

but it didn't work out, it got ugly lol

OpenStudy (here_to_help15):

Unusual question! Let's go with Tim first. Suppose Tim encounters the radical equation \[\sqrt{x} + 9=0\] Tim tries to solve this by moving the 9 to the other side of the equation; he subtracts 9 from both sides and obtains \[\sqrt{x} = - 9\] To eliminate the radical and to solve for x, he now squares both sides of this equation. Can you help him do that? What possible solution do the two of you arrive at? Substitute this possible solution back into the original equation. Is the resulting equation true or false? How about you try supervising John. Guide him in the right direction to create and solve a radical equation whose solution really is a solution, that is, it satisfies the original equation.

OpenStudy (anonymous):

You're copying that right off someone else on here. I've read that one already. Thanks for trying though.

OpenStudy (here_to_help15):

As you know im quoting this since i have very few knowledge on what this question is asking i will provide link http://openstudy.com/study#/updates/52c5ce14e4b0b729fb8cc9fe

OpenStudy (anonymous):

I know, but it didn't help me lol

OpenStudy (here_to_help15):

Yea :) Sorry i should have said where i was getting it from first

OpenStudy (anonymous):

No, it's okay . Radicals aren't my strong suit lol i'll figure it out, but thank you for trying!